To a crude first approximation, an electron in a linear polyene may be considered to be a particle in a one-dimensional box. The polyene - carotene contains 22 conjugated C atoms, and the average inter-nuclear distance is 140 pm. Each state up to n =11 is occupied by two electrons. Calculate the separation in energy between the ground state and the first excited state in which one electron occupies the state with n=12.
Answer #1

Particle (electron ) in a one diamentional box ,Energy equation is,

En = n2 h2 /8mL2

Where n is the quantum number, 1,2,3 ...

m isthe mass of the particle (electron)

L is the length of the barrior

Here En is n=11 to n=12

so the change in energy ΔE=E12-E11

ΔE=(122-112)*h2/8mL2

(144-122)h2/8mL2

=23h2/8mL2

Here the average internuclear distance L=140 pm=1.4*10-10m

In polyene - carotene contains 22 conjugated C atoms

so we want tocalculat 22 conjugateted c atoms L =22*140pm =3080pm=3.08*10-9m

ΔE12-->ΔE11= 23h2/8mL2

=(23*(6.626*10-34)2)/(8*(9.1*10-31Kg)*(3.08*10-9)2)

= 1009.789*10-68/690.60*10-49

=1.463*10-19J